General Physics Data Sheet Diameter of bob31m Radius of boba--..no Length of m Number of Pendulem I m) Vibrations Length of Total Time()Period Ts) Squareof Strine () Period T 83 1 12.44 4 LS Slope of graph =- Acceleration Due to Gravity (m/s) %Error Standard


how would I figure out the slope of the graph, standard acceleration, experimental, and percent error? also how would I graph this?
Answer #1

Drew a graph between T^2(y axis) and m(x axis)

slope of the graph

gradient=m=tan\theta=dy/dx=\Delta y/\Delta x=slope

gradient=\frac{y_2-y_1}{x_2-x_1}

standard acceleration

as we know

T=2\pi \sqrt{\frac{L}{g}}

T^2=4\pi ^2{\frac{L}{g}}

g=4\pi ^2{\frac{L}{T^2}}

experimental acceleration

it is the slope of the graph

percent error

\frac{ExperimentalValue-Theoretical Value}{Theoretical value}\times 100   %

here theoretical value of g is   9.81 m/s^2